OpenStudy (anonymous):

hoose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of one half. y − 6 = one half(x + 3) y = one halfx − 6 y + 3 = one half(x − 6) x − 2y = 12

jimthompson5910 (jim_thompson5910):

point slope form y-y1 = m(x-x1) y-y1 = 1/2(x-x1) ... plug in m = 1/2 y-(-3) = 1/2(x-6) ... plug in (6,-3) for x1 and y1 y+3 = 1/2(x-6)

OpenStudy (anonymous):

so witch ones the answer

jimthompson5910 (jim_thompson5910):

y+3 = 1/2(x-6) is the same as y+3 = one half(x-6)

jimthompson5910 (jim_thompson5910):

since 1/2 = one half

OpenStudy (anonymous):

oh i got it thank you so much

jimthompson5910 (jim_thompson5910):

np